A 14.0-m uniform ladder weighing 520 N rests again
Answer #1

= 830 (4/14) cos 55 = 136.02 N

b)

force towards the wall = Static frictional force

830 ( 9.4/14) cos 55 = us x 1350

Us = 0.236